A gaseous mixture contains 415.0 tour H2 (g), Torr N2 (g), and 92.1 Torr Ar (g). Caluclate the mole fraction, x, of each of these gases. XH2=?, XN2=?, XAr=?
Answer #1

lets,

We can express the mole function Xi  of an individual gas i in a mixture in terms of the component's

partial pressure Pi

X_{i}=\frac{P_{i}}{P_{tota}} .........................................(1)

so, P_{i}=X_{i}*P_{total}...............................(2)

the value of  P^{_{H_{2}}}=415.0 tour

P_{N_{2}}=415.0 tour

P_{Ar}=92.1 tour   

P_{total}=415.0 +415.0+92.1 tour=922.1 tour

  X_{H_{2}}=\frac{415.0}{922.1}=0.4500

  X_{N_{2}}=\frac{415.0}{922.1}=0.4500

  X_{Ar}=\frac{92.1}{922.1}=0.0999

H2 mole fraction =.4500

N2 mole fraction =.4500

Ar mole fraction is =.0999